OpenStudy (anonymous):

Ronald drives along a mountain and enters a tunnel with a speed of 30m/s, suddenly he realizes that 155m in front of him is a truck going at 5m/s. Ronald's acceleration is 2m/s^2. how long will it take before he hits the car?

OpenStudy (anonymous):

@rajathsbhat @hartnn @ajprincess @experimentX

OpenStudy (anonymous):

these are the things that i know:

OpenStudy (anonymous):

a= 2m/s^2 s=155m relative v= 25m/s

OpenStudy (anonymous):

the question might not be grammatically correct because i had to translate from Dutch.. :/

OpenStudy (anonymous):

155 + 30t + t^2 = 5t t^2 +25t + 155 =0 Solve for t

OpenStudy (anonymous):

ah thx!

OpenStudy (anonymous):

but why is it 30?

OpenStudy (anonymous):

Distance covered by ronald = Distance Covered By Truck \[S = ut + \frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

sorry, i didn't see that it was 25..

OpenStudy (anonymous):

S = u*t

OpenStudy (anonymous):

truck isn't accl right?

OpenStudy (anonymous):

yes :) i got the answer 11,4s